OpenStudy (anonymous):

Combine like terms and simplify : 2 + 4x - y - 2x + 7 + 3y

OpenStudy (ivettef365):

do you know what are like terms ?

OpenStudy (anonymous):

2 x+2 y+9.

OpenStudy (anonymous):

no what are they

OpenStudy (ivettef365):

Like terms" are terms whose variables (and their exponents such as the 2 in x2) are the same. In other words, terms that are "like" each other.

OpenStudy (ivettef365):

which terms look alike

OpenStudy (anonymous):

im actually glad u are trying to help instead of telling me the answers

OpenStudy (ivettef365):

you can combine the 2 and the 7 2 + 7 = also, can combine 4x and -2x, so you have 4x - 2x = ??

OpenStudy (anonymous):

Is receiving the answer not helping?

OpenStudy (anonymous):

it is and im actually glad u gave it to me but im even happier ivettef is tryning to help learn it

OpenStudy (ivettef365):

do you understand how is done?

OpenStudy (anonymous):

no not yet im still not understating

OpenStudy (anonymous):

how do you combine like terms

OpenStudy (ivettef365):

ok, 2 + 7 you know that one = 9 then, 4x - 2x (the variable does not change, only the numbers) so that is (4-2) = 2 so it is 2x then you have -y and 3y, you can put as 3y - y, which is (3-1) = 2y so you have 2x+2y+9